88). Perimeter of a rectangle. The perimeter of a rectangular
backyard is 6x + 6 yards. If the width is x yards, find a
binomial that represents the length.

Answer by nerdybill(7384)

88). Perimeter of a rectangle. The perimeter of a rectangular
backyard is 6x + 6 yards. If the width is x yards, find a
binomial that represents the length.
.
We know:
perimeter = 2L + 2W
where
L = length
W = width
.
Plug in what was given and solve for L:
perimeter = 2L + 2W
6x + 6 = 2L + 2x
Dividing both sides by 2:
3x + 3 = L + x
Subtracting both sides by x:
2x + 3 = L
2x + 3 Yards (this is your binomial expression for length)